OpenStudy (anonymous):

Guys, a simple question please How can I turn " M=e^t - e^tM " into " M= e^t / (1 + e^t) "

OpenStudy (amistre64):

equate them and see what you can do to them :) maybe

OpenStudy (anonymous):

m+e^tm=e^t m(1+e^t)=e^t m=e^t/(1+e^t)

OpenStudy (amistre64):

e^t -e^(tM) = e^t/(1+e^t) ; *(1+e^t) both sides (e^t -e^(tM))(1+e^t) = e^t

OpenStudy (amistre64):

e^t +e^2t - e^(tM) - e^(t^2M) = e^t

OpenStudy (anonymous):

Oh, right, I've got it! Thank you very much ! :D